Solve: 2x^2-3x-5=0
Please show me the steps to get the answer. Thanks

Answer by edjones(8007)

2x^2-3x-5=0
2*-5=-10; two factors of -10 whose sum is -3 are -5 and +2.
2x^2+2x-5x-5=0
2x(x+1)-5(x+1)=0 Factoring by grouping.
(2x-5)(x+1)=0
x=-1, x=5/2
.
Ed